OpenStudy (anonymous):

y-8=-x please graph on a table and tell why you did what you did.

OpenStudy (anonymous):

First let's rearrange this to the form y=mx+b we get y=-x+8 Next, we look at when x =0 to give us our y-intercept y =-(0) +8 --> y=8, when x =0 then we look for our roots, also known as x intercept 0 = -x +8 --> x =8 , when y = 0. This is our x intercept or root. Let's just plot this for now.  Now we know that the slop is defined as m, which in this case is -1. This means that as we run--- 1 unit over on x axis, we rise -1 unit on the y axis. Overall our graph will look something like this  Do you have anymore questions regarding this answer?